please show how you solved this. thank you!ООО Q Search Sheet Chapter_7_Inc-Class Group Work 1 (1) View Formulas Data Review + Share Home K16 Insert x Page Layout fx SS
Answer #1

a)

Actual results Flexible budget variance (Actual -flexible) Flexible budget (Mater budget -flexible budget) Master Budget
sales 46000*60= 2760000 230000 F 46000*55= 2530000 55000 F 45000*55= 2475000
variable cost 46000*34= 1564000 46000 F 46000*35= 1610000 35000 U 45000*35= 1575000
contribution margin 1196000 276000 F 920000 20000 F 900000
Fixed cost 627000 33000 F 660000 0 NA 660000
Net income 569000 309000 F 260000 20000 F 240000

**If Actual revenue is higher than Flexible budget revenue then variance is favorable (and vise versa)

If actual cost is less than flexible budget then variance is favorable (and vise versa)

**If Flexible budget revenue is higher than master budget revenue ,then variance is favorable (and vise versa)

If cost is less than master budget then variance is favorable (and vise versa)

2)Actual amount spent for material B= Flexible budget variance + unfavorable variance -Favorable variance

        = 39000 + 800 -400

        = 39400